π1= -Q²+17Q-42 π2 = -Q²+16Q-38, at what level of output will profit be zero?


1
Expert's answer
2021-12-03T13:04:31-0500

Given; π1= -Q²+17Q-42 π2 = -Q²+16Q-38


The output will zero when "\frac{\delta\pi}{\delta Q}=0"


For π1= -Q²+17Q-42


"\frac{\delta\pi}{\delta Q}=-2Q+17"


"Q=\frac{17}{2}=8.5"


For π2 = -Q²+16Q-38;

"\frac{\delta\pi}{\delta Q}=-2Q+16"


"Q=\frac{16}{2}=8.0"